OpenStudy (freckles):

Hey! I want to think about evaluating: \[\int_0^\infty \lfloor x\rfloor e^x dx \text{ where } \lfloor . \rfloor\text{ is greatest integer function }\]

OpenStudy (freckles):

trying to figure out the floor function thingy

OpenStudy (freckles):

got it that is a long command for the floor function

OpenStudy (freckles):

anyways... \[n \le x<n+1 \text{ then } \lfloor x \rfloor=n \text{ where } n \text{ is integer }\] now I think I want to do integration by parts...

OpenStudy (freckles):

trying to think how to integrate the greatest integer function

OpenStudy (freckles):

\[\int\limits_0^\infty \lfloor x \rfloor e^x dx \\ \int\limits_0^1 0 \cdot e^x dx+\int\limits_1^2 1 \cdot e^x + \int\limits_2^3 2 \cdot e^x + \cdots + \int\limits_n^{n+1} n \cdot e^{x} dx +\cdots \]

OpenStudy (freckles):

\[\sum_{n=1}^{\infty} \int\limits _n^{n+1} n \cdot e^x dx\]

OpenStudy (freckles):

\[\sum_{n =1}^\infty n \int\limits _n^{n+1} e^x dx \\ \sum_{n=1}^\infty n e^x|_n^{n+1} \\ \sum_{n=1}^\infty n[e^{n+1}-e^{n}]\]

OpenStudy (freckles):

actually I guess I don't need integration by parts

ganeshie8 (ganeshie8):

thats looks very neat, but it doesn't converge right

OpenStudy (freckles):

yeah wolfram says it doesn't diverge I thought this was suppose to converge those for some reason did i make a mistake above

OpenStudy (freckles):

I mean it does diverge sorry

OpenStudy (freckles):

you know what maybe if the exponent on the x was negative it would have converged

ganeshie8 (ganeshie8):

yeah exponential overtakes any polynomial, floor(x) is cheaper than a linear polynomial

OpenStudy (freckles):

did something wrong

OpenStudy (freckles):

oops didn't mean to delete that without copying it

OpenStudy (freckles):

\[\sum_{n=1}^{\infty}n(e^{-(n+1)}-e^{-n})\] whatever I know this converges I just need to find the sum now I was thinking this maybe a telescoping series

ganeshie8 (ganeshie8):

\[\int\limits_{0}^{\infty} \lfloor x \rfloor e^{-x} dx \\ =\sum_{n=1}^{\infty} \int\limits_{n}^{n+1} n e^{-x} dx \\ =\sum_{n=1}^{\infty} -n e ^{-x}|_n^{n+1} \\ \\ =\sum_{n=1}^{\infty} -n[e^{-(n+1)}-e^{-n}] \\ \text{ ratio test: } \\ \lim_{n \rightarrow \infty} \frac{-(n+1)[e^{-(n+2)}-e^{-(n+1)]}}{-n [e^{-(n+1)}-e^{-n}]}\] \[\lim_{n \rightarrow \infty}-(1+\frac{1}{n}) \frac{ e^{-(n+2)+(n+1)-e^{-(n+1)+(n+1)}}}{e^{-(n+1)+(n+1)}-e^{-n+(n+1)} } \\ -(1+0) \frac{e^{-1}-e^{0}}{e^{0}-e^{1}} =-1 \frac{e^{-1}-1}{1-e} =-1 \frac{1-e}{e-e^2} =\frac{e-1}{e-e^2} \\ =\frac{e-1}{e(1-e)}=\frac{-1}{e}\]

OpenStudy (freckles):

\[(e^{-2}-e^{-1})+2(e^{-3}-e^{-2})+3(e^{-4}-e^{-3})+4(e^{-5}-e^{-4}) +\cdots \] not exactly a telescoping series the way it is written at last

ganeshie8 (ganeshie8):

i pulle that from ur deleted replies

OpenStudy (freckles):

thanks @ganeshie8 now people can know what problem I'm talking about I'm trying to evaluate: \[\int\limits_{0}^{\infty} \lfloor x \rfloor e^{-x} dx\]

ganeshie8 (ganeshie8):

mimicing ur earlier work, \[\int\limits_{0}^{\infty} \lfloor x \rfloor e^{-x} dx = (e-1)\sum\limits_{n=1}^{\infty} \frac{n}{e^n} \]

OpenStudy (xapproachesinfinity):

so the first floor (x) e^x diverge or not?

OpenStudy (freckles):

i showed it diverge so I think it diverge (unless I made some mistake)

OpenStudy (freckles):

well wolfram showed my sum diverged :p

OpenStudy (xapproachesinfinity):

\(\color{blue}{\text{Originally Posted by}}\) @freckles \[\sum_{n =1}^\infty n \int\limits _n^{n+1} e^x dx \\ \sum_{n=1}^\infty n e^x|_n^{n+1} \\ \sum_{n=1}^\infty n[e^{n+1}-e^{n}]\] \(\color{blue}{\text{End of Quote}}\) from tis result ?

OpenStudy (freckles):

I could have use the ratio test to come up with that know

OpenStudy (freckles):

ues from that last line

OpenStudy (xapproachesinfinity):

ah ok i guess ration test can testify divergence

ganeshie8 (ganeshie8):

we can simply use the limit test

OpenStudy (xapproachesinfinity):

oh yes! that limit does not go to 0

ganeshie8 (ganeshie8):

\[\sum\limits_{n=1}^{\infty} \frac{n}{e^{nx}}=\sum\limits_{n=1}^{\infty} \frac{d}{dx}e^{-nx}=\frac{d}{dx}\sum\limits_{n=1}^{\infty} e^{-nx}= \frac{d}{dx}\frac{e^{-x}}{1-e^{-x}}=\cdots\]

OpenStudy (freckles):

and we know \[\frac{d}{dx} \frac{e^{-x}}{1-e^{-x}} \\ \frac{d}{dx} \frac{1}{e^x-1} \\ \frac{d}{dx}(e^x-1)^{-1} \\ -1(e^x-1)^{-2}(e^x) \\ - \frac{e^x}{(e^x-1)^2}\] but we should evaluate this at x=1 since we had that we really wanted to evaluate \[\int\limits\limits_{0}^{\infty} \lfloor x \rfloor e^{-x} dx = (e-1)\sum\limits_{n=1}^{\infty} \frac{n}{e^n} \] \[(e-1) \frac{-e}{(e-1)^2}=\frac{-e}{e-1}\] but I think we are to get 1/(e-1)

ganeshie8 (ganeshie8):

Oh right \[\int\limits\limits_{0}^{\infty} \lfloor x \rfloor e^{-x} dx =-\sum_{n=1}^\infty n[e^{-(n+1)}-e^{-n}] = \frac{e-1}{e}\sum\limits_{n=1}^{\infty} \frac{n}{e^n}\]

ganeshie8 (ganeshie8):

algebra mistakes

OpenStudy (freckles):

thanks @ganeshie8 ok I was sorta failing at the algebra a little too but I see it now

OpenStudy (freckles):

I can't remember who gave me this problem but @Loser66 this is also on your gre practice exam if you want to look at this.

ganeshie8 (ganeshie8):

i remember working this problem differently sometime back here in openstudy but google isn't helping at the moment..

OpenStudy (freckles):

And I bet you have answered so many questions since this that going back through your old questions is horrifying and it also doesn't work to well sometimes. Sometimes it automatically throws me out of the looking through my old question thing. Not sure if you know what I mean.

OpenStudy (loser66):

Thanks so mmmmmmmmuch

OpenStudy (xapproachesinfinity):

truly i could not see how you threw differentiation in there @gabylovesu how is that sum equal to that differentiation ?

OpenStudy (xapproachesinfinity):

@ganeshie8

OpenStudy (xapproachesinfinity):

wrong tagging lol

OpenStudy (anonymous):

Have you tried parameterizing and applying the Laplace transform? I'm thinking something along the lines of \[I(s)=\int_0^\infty \lfloor x\rfloor e^{-sx}\,dx=\mathcal{L}\{\lfloor x\rfloor\}\] The transform definitely exists because \(\lfloor x\rfloor\) is piecewise continuous and of exponential order.

ganeshie8 (ganeshie8):

\[\color{red}{\sum\limits_{n=1}^{\infty} \frac{n}{e^{nx}}=-\sum\limits_{n=1}^{\infty} \frac{d}{dx}e^{-nx}}=-\frac{d}{dx}\sum\limits_{n=1}^{\infty} e^{-nx}= -\frac{d}{dx}\frac{e^{-x}}{1-e^{-x}}=\cdots\] I think it would be easier to make sense of it by differentiating \(\large \color{red}{e^{-nx}}\) and seeing that you get back the starting expression...

OpenStudy (freckles):

hey @SithsAndGiggles I haven't thought of that but I would be interested in seeing that way if you wanted to show that way. I honestly I haven't seen any laplace transform action in like 10 years so I kind of forgot all of that. :p

OpenStudy (anonymous):

The idea would be to express \(\lfloor x\rfloor e^{-x}\) in terms of the step function \[\theta(x-c)=\begin{cases}1&\text{for }x\ge c\\0&\text{for }x<c\end{cases}\]

OpenStudy (xapproachesinfinity):

ooh i see it now

OpenStudy (anonymous):

\[\lfloor x\rfloor e^{-x}=\sum_{c=1}^\infty ce^{-x}\Bigg(\theta(x-c)-\theta(x-c-1)\Bigg)\] So \[\begin{align*}\mathcal{L}\left\{\lfloor x\rfloor\right\}&=\int_0^\infty \lfloor x\rfloor e^{-sx}\,dx\\ &=\int_0^\infty \sum_{c=1}^\infty ce^{-sx}\Bigg(\theta(x-c)-\theta(x-c-1)\Bigg)\,dx\\ &=\sum_{c=1}^\infty c\int_0^\infty e^{-sx}\Bigg(\theta(x-c)-\theta(x-c-1)\Bigg)\,dx\\ &=\sum_{c=1}^\infty c\mathcal{L}\{\theta(x-c)-\theta(x-c-1)\}\\ &=\sum_{c=1}^\infty c\frac{e^{-cs}-e^{-(c+1)s}}{s} \end{align*}\] Setting \(s=1\) gives the series @ganeshie8 was working with. Nothing new I suppose :P

OpenStudy (freckles):

does seem similar but still pretty @SithsAndGiggles anyways this is making me hungry so peace and thanks for the fun

OpenStudy (freckles):

I will come back later though to analyze more deeply what you said

OpenStudy (anonymous):

As for the actual result: \[\sum_{c=1}^\infty c(e^{-c}-e^{-(c+1)})=\sum_{c=1}^\infty c\left(\frac{1}{e^c}-\frac{1}{e^{c+1}}\right)=\frac{1}{e}+\frac{1}{e^2}+\frac{1}{e^3}+\cdots=\frac{1}{1-\frac{1}{e}}-1\]